A proton (+q =+1.602 x 10^-19) moves along a straight line from point a to point b with a separation distance d = 0.50m. Considering the electric field along this line uniform magnitude of 1.50 x 10^7V/m and directed from point a to point b.



Determine:


a.) the force on the proton



b.) the work done on it by the field (in Joules & eV units)



c.) the potential difference (V ab)

1
Expert's answer
2022-03-10T11:15:39-0500

a.) the force on the proton

"F=qE\\
=1.602*10^{-19}*1.50*10^7=2.4*10^{-12}\:\rm N"


b.) the work done on it by the field (in Joules & eV units)

"W=Fd\\
=2.4*10^{-12}*0.50=1.2*10^{-12}\:\rm J=7.5*10^6\: eV"

c.) the potential difference (V ab)

"V_{ab}=Ed\\
=1.50*10^7*0.50=7.5*10^6\:\rm V"